x^2-x-6 over x^2-2x-8 Divided by x^2-9 over x^2+7x+12



I had no clue how to write this equation out


Found 2 solutions by edjones, sowmya:
Answer by edjones(7311)

It's not an equation.
(x^2-x-6)/(x^2-2x-8) / (x^2-9)/(x^2+7x+12)
=[(x-3)(x+2)]/[(x-4)(x-2)] * [(x+4)(x+3)]/[(x+3)(x-3)] Factor and put in a times sign and invert the 2nd fraction.
=[(x+2)]/[(x-4)(x-2)] * (x+4) Cancel.
=(x^2+6x+8)/(x^2-6x+8)
.
Ed

Answer by sowmya(32)

x^2-x-6 over x^2-2x-8 Divided by x^2-9 over x^2+7x+12

solution:
Numerator = x^2-x-6 / x^2-2x-8
=x^2-x-6 /(x-4)(x+2)
=x-3)(x+2)/(x-4)(x+2)=x-3/x+4
denominator = (x-3)(x+3)/(x+4)(x+3)= (x-3)/x+4
nm/dn = 1
(QED)